# On Bayesian Neural Networks with Dependent and Possibly Heavy-Tailed Weights

### v4

@@ 7.4 Simulation of Model 2
Moreover, for $\ell=0,\ldots,L$ and $r,s=1,\ldots,d$, we denote by $H_{rs}^{(\ell)}(\mathbb{x})$ the $rs$-entry of $\mathbb{H}^{(\ell)}(\mathbb{x})$ and by $H_{rs}^{(\ell),\sharp}(\mathbb{x})$ the $rs$-entry of $\mathbb{H}^{(\ell)}(\mathbb{x})]^{\sharp}$ (i.e., the square-root of $\mathbb{H}^{(\ell)}(\mathbb{x})$). Note that, given $\mathbb{H}^{(\ell-1)}(\mathbb{x})$, $\{\gamma_{j}^{(\ell-1)}(\mathbb{x})\}_{j=1,\ldots,n_{\ell}}$ are independent and identically distributed with law $\mathcal{N}_{d}(0,\mathbb{H}^{(\ell-1)}(\mathbb{x}))$. Note also that
@@
$\mathbb{H}^{(\ell)}(\mathbb{x})=C_{B}1_{d}1_{d}^{\top}+C_{W}\sigma\left(\widetilde{\mathbb{N}}^{(\ell)}[\mathbb{H}^{(\ell-1)}(\mathbb{x})]^{\sharp}\right)^{\top}\mathbb{V}^{(\ell)}\sigma\left(\widetilde{\mathbb{N}}^{(\ell)}[\mathbb{H}^{(\ell-1)}(\mathbb{x})]^{\sharp}\right)$ (38)
$\boldsymbol{\gamma}^{(\ell-1)}(\mathbb{x})=\left(\widetilde{\mathbb{N}}^{(\ell)}[\mathbb{H}^{(\ell-1)}(\mathbb{x})]^{\sharp}\right)^{\top}.$ (39)
@@
Lemma 7.1.
Under the foregoing assumptions and notation, it holds
$\mathbb{Z}^{(L+1)}(\mathbb{x})\overset{\mathcal{L}}{=}\widetilde{\mathbb{N}}^{(L+1)}[\mathbb{H}^{(L)}(\mathbb{x})]^{\sharp}.$
Hence
$\phi(\boldsymbol{\theta};\mathbb{Z}^{(L+1)}(\mathbb{x}))=\mathbb{E}\left[\exp\left(-\frac{1}{2}\mathrm{Tr}\big[\boldsymbol{\theta}^{\top}\mathbb{H}^{(L)}(\mathbb{x})\boldsymbol{\theta}\big]\right)\right],\quad\boldsymbol{\theta}\in\mathbb{R}^{d\times n_{L+1}}.$ (40)
Lemma 7.2.
Let $\{V_{n,j}\}_{n\geq 1,j=1,\ldots,n}$ be a sequence of non-negative random variables such that
$\sum_{j=1}^{n}V_{n,j}\overset{\mathcal{L}}{\to}{\rm ID}(a,\rho),\quad\text{as $n\to\infty$}$
where $a\geq 0$ is a non-negative constant and $\rho$ is a Lévy measure on $(0,\infty)$, and let $\{\xi_{j}\}_{j\geq 1}$ be a sequence of independent random vectors with law $\mathcal{N}_{d}(0,\mathbb{C})$, independent of $\{V_{n,j}\}_{n\geq 1,j=1,\ldots,n}$. If $\sigma$ satisfies the assumption $(i)$ of Theorem 4.1, then
@@
Proof of Theorem 4.2.
For ease of notation, we omit the explicit reference to the input matrix $\mathbb{x}$ in every symbol it occurs. Since vec and $\top$ are isometries of Euclidean spaces, we equivalently prove the theorem for the random matrices $\mathbb{Z}^{(L+1)}$ and $\mathbb{G}^{(L+1)}$ rather that for the random vectors $Z^{(L+1)}$ ad $G^{(L+1)}$. Let $\mathcal{S}_{+}^{d}$ be the space of the order $d$ symmetric and positive semi-definite real matrices endowed with the Frobenius norm, and note that, for any fixed $\boldsymbol{\theta}\in\mathbb{R}^{d\times d}$, the mapping $\mathcal{S}_{+}^{d}\ni\mathbb{A}\mapsto\exp\left(-\frac{1}{2}\mathrm{Tr}\big[\boldsymbol{\theta}^{\top}\mathbb{A}\boldsymbol{\theta}\big]\right)$ is real-valued, continuous and bounded. Hence, by (40) and the definition of weak convergence it holds that,
$\mathbb{H}^{(L)}\xhookrightarrow{\pi\mathcal{L}}\mathbb{K}^{(L+1)}\Longrightarrow\pi\overrightarrow{\lim}\phi(\boldsymbol{\theta};\mathbb{Z}^{(L+1)})=\mathbb{E}\left[\exp\left(-\frac{1}{2}\mathrm{Tr}\big[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}\boldsymbol{\theta}\big]\right)\right]=\phi(\boldsymbol{\theta};\mathbb{G}^{(L+1)}).$
Therefore to prove the theorem it suffices to prove that $\mathbb{H}^{(L)}\xhookrightarrow{\pi\mathcal{L}}\mathbb{K}^{(L+1)}$, for every permutation $\pi$ of $[L]$. We proceed by induction on the number of layers $L$, the basis of the induction being $\mathbb{H}^{(1)}\xhookrightarrow{\pi\mathcal{L}}\mathbb{K}^{(2)}$ which reduces to
$\mathbb{H}^{(1)}{\xrightarrow[n_{1}\to\infty]{\mathcal{L}}}\mathbb{K}^{(2)}$ (41)
because, for $L=1$, the unique permutation of $[L]$ is the identity. We defer the proof of (41) after the proof of the inductive step. So, let $L\geq 3$ and assume
$\mathbb{H}^{(L-1)}\xhookrightarrow{\tau\mathcal{L}}\mathbb{K}^{(L)}\quad\text{for every permutation $\tau$ of $[L-1]$}.$ (42)
@@
$\pi\lim\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})=\phi(\boldsymbol{\theta};\mathbb{K}^{(L+1)})\quad\text{$\forall\boldsymbol{\theta}\in\mathbb{R}^{d\times d}$ and every permutation $\pi$ of $[L]$},$ (43)
assuming (41) (to be proved) and (42). Let $\pi$ be any permutation of $[L]$ and let $j^{*}=\pi^{-1}(L)$. Hence $(\pi\preceq j^{*}-1)=(\tau\preceq j^{*}-1)$ for some permutation $\tau$ of [L1][L-1]. Since the mapping $\mathbb{A}\to\mathbb{A}^{\sharp}$ is continuous on $\mathcal{S}_{+}^{d}$, the inductive step (42) yields $\mathbb{H}^{(L-1)}]^{\sharp}\xhookrightarrow{(\pi\preceq j^{*}-1)\mathcal{L}}\mathbb{K}^{(L)}_{(\pi\preceq j^{*}-1)}$. Since, by construction, $\mathbb{H}^{(L-1)}$ is independent of $\mathbb{V}^{(L)}$ and $\widetilde{\mathbb{N}}^{(L)}$ and $\sigma$ is continuous, by (38) it follows that
$\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{H}^{(L-1)}]^{\sharp}\right)^{\top}\!\mathbb{V}^{(L)}\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{H}^{(\ell-1)}]^{\sharp}\right)\xhookrightarrow{(\pi\preceq j^{*}\!-\!1)\mathcal{L}}\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}\right)^{\top}\mathbb{V}^{(L)}\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}\right).$
Note that also $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}-1)}]^{\sharp}$ is independent of $\mathbb{V}^{(L)}$ and $\widetilde{\mathbb{N}}^{(L)}$. Therefore, if $\mathbb{\Xi}:=[\xi_{1},\ldots,\xi_{n_{L}}]\in\mathbb{R}^{d\times n}$ is the matrix defined by $\mathbb{\Xi}=\widetilde{\mathbb{N}}^{(L)}[\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}$, then $\mathbb{\Xi}$ is independent of $\mathbb{V}^{(L)}$ and, given $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}$, $\{\xi_{1},\ldots,\xi_{n_{L}}\}$ are independent and identically distributed random vectors in $\mathbb{R}^{d}$ with law $\mathcal{N}_{d}(0,\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)})$ each. Consequently, in view of (37), one has
$(\pi\preceq\!j^{*}\!-\!1)\overrightarrow{\lim}\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})=\mathbb{E}\left[\exp\left(\mathbb{i}\langle\boldsymbol{\theta},C_{B}\mathbb{1}_{d}\mathbb{1}_{d}^{\top}+C_{W}\sum_{j=1}^{n_{L}}V_{n_{L}j}^{(L)}\sigma(\xi_{j})\sigma(\xi_{j})^{\top}\rangle_{F}\right)\right].$
On the other hand, with $n_{\pi(j^{*})}=n_{L}$, by Lemma 7.2 we have that
@@
where, referring to (6), $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}$ is defined similarly to $\mathbb{K}^{(L+1)}$ but with the $\xi_{j}$’s in place of the $\zeta_{j}^{(L)}$’s and $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}$ in place of $\mathbb{K}^{(L)}$, i.e.
$\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}:=C_{B}1_{d}1_{d}^{\top}+C_{W}\Biggl(a^{(L)}\mathbb{E}[\sigma(\xi_{1})\sigma(\xi_{1})^{\top}\,|\,\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]+\sum_{j=1}^{N_{L}((0,\infty))}T_{j}^{(L)}\sigma(\xi_{j})\sigma(\xi_{j})^{\top}\Biggr).$ (44)
Since $L=\pi(j^{*})$ we have proved that
$(\pi\preceq j^{*})\overrightarrow{\lim}\phi(\boldsymbol{\theta};\mathbb{H}^{L})=\lim_{n_{L}\to\infty}\left[(\pi\preceq\!j^{*}\!-\!1)\overrightarrow{\lim}\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})\right]=\mathbb{E}\left[\exp\left(\mathbb{i}\mathrm{Tr}\left[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}\right]\right)\right].$ (45)
If $j^{*}=L$, then we are done. Indeed, since $(\pi\preceq j^{*})=\pi$ and $(\pi\preceq j^{*}\!-\!1)=\tau$ for some permutation $\tau$ of [L1][L-1], it follows that $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}=\mathbb{K}^{(L)}$ by (42). Hence $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}=\mathbb{K}^{(L+1)}$ by (44), after recalling (6). We thus assume $j^{*}\not=L$ and prove that $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}\xhookrightarrow{(\pi\succeq j^{*}+1)\mathcal{L}}\mathbb{K}^{(L+1)}$. Since $(\pi\succeq j^{*}\!+\!1)=(\tau\succeq j^{*}\!+\!1)$ for some permutation $\tau$ of $[L]$, it follows that $\mathbb{K}^{(L)}_{(\pi\preceq j^{*})}\xhookrightarrow{(\pi\succeq j^{*}+1)\mathcal{L}}\mathbb{K}^{(L)}$ by (42) . A simple computation with characteristic functions shows
$(\xi_{1},\ldots,\xi_{n_{L}})\xhookrightarrow{(\pi\succeq j^{*}+1)\mathcal{L}}(\zeta_{1}^{(L)},\ldots,\zeta_{n_{L}}^{(L)}).$ (46)
@@
$\mathbb{E}|X_{n}(i_{1})|^{p_{1}}\ldots|X_{n}(i_{k})|^{p_{k}}\to\mathbb{E}|X(i_{1})|^{p_{1}}\ldots|X(i_{k})|^{p_{k}}.$
Since $\sigma$ is continuous and is either bounded or with subpolynomial growth, using the dominated convergence theorem in the bounded case and the generalized dominated convergence theorem in the case of a subpolynomial growth, one has that
$\tilde{\mathbb{E}}\left[\sigma(\xi_{1})\sigma(\xi_{1})^{\top}\,|\,\mathbb{K}^{(L)}_{(\pi\preceq j^{*})}\right]\xhookrightarrow{(\pi\succeq j^{*}+1)\tilde{\mathbb{P}}}\tilde{\mathbb{E}}\left[\sigma(\zeta_{1}^{(L)}\sigma(\zeta_{1}^{(L)})^{\top}\,|\,\mathbb{K}^{(L)}\right].$
Using the continuity of $\sigma$, one also has that
$\sum_{j=1}^{N_{L}((0,\infty))}T_{j}^{(L)}\sigma(\xi_{j})\sigma(\xi_{j})^{\top}\xhookrightarrow{(\pi\succeq j^{*}+1)\tilde{\mathbb{P}}}\sum_{j=1}^{N_{L}((0,\infty))}T_{j}^{(L)}\sigma(\zeta_{j}^{(L)})\sigma(\zeta_{j}^{(L)})^{\top}$
where, with a little abuse of notation we still denote by $N_{L}=\{T_{j}^{(L)}\}$ a copy of $N_{L}=\{T_{j}^{(L)}\}$ on $\tilde{\Omega}$. This convergence is clear if $\rho^{(L)}((0,\infty))<\infty$. In the case when $\sigma$ is bounded this convergence holds by the dominated convergence theorem. Indeed, being $\rho^{(L)}$ a Lévy measure, one has that the versions on $\tilde{\Omega}$ of the random variables
$\sum_{j=1}^{N_{L}((0,1))}T_{j}^{(L)}\quad\text{and}\quad\sum_{j:\,T_{j}^{(L)}\in[1,\infty)}T_{j}^{(L)}$
@@
$\begin{split}\pi\overrightarrow{\lim}\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})&=(\pi\succeq j^{*}+1)\overrightarrow{\lim}\left[(\pi\preceq j^{*})\overrightarrow{\lim}\phi(\boldsymbol{\theta};\mathbb{H}^{L})\right]\\ &=(\pi\succeq j^{*}+1)\overrightarrow{\lim}\mathbb{E}\left[\exp\left(\mathbb{i}\mathrm{Tr}\left[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}\right]\right)\right]\\ &=\mathbb{E}\left[\exp\left(\mathbb{i}\mathrm{Tr}\left[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}\right]\right)\right]\end{split}$
and (43) is proved. It remains to prove (41). By construction $\mathbb{H}^{(0)}=\mathbb{K}^{(1)}$. Hence, with $\ell=L=1$, by (39), for $j=1\ldots,n_{1}$, it holds that $\gamma_{j}^{(0)}\overset{\mathcal{L}}{=}\zeta_{j}^{(1)}$, where $\zeta_{j}^{(1)}$ has law $\mathcal{N}_{d}(0,\mathbb{K}^{(1)})$. Consequently, by (37), one has
@@
Hence, after recalling (6), by Lemma 7.2, $\mathbb{K}^{(2)}$ is the limit in law of $\mathbb{H}^{(1)}$ as $n_{1}\to\infty$. The proof is thus completed. ∎
Acknowledgments.
The authors acknowledge the support of the Istituto Nazionale di Alta Matematica - gnampa group.

### v5

@@ Appendix A Proof of Theorem 2
Moreover, for $\ell=0,\ldots,L$ and $r,s\in[d]$, we denote by $H_{rs}^{(\ell)}(\mathbb{x})$ the $rs$-entry of $\mathbb{H}^{(\ell)}(\mathbb{x})$ and by $H_{rs}^{(\ell),\sharp}(\mathbb{x})$ the $rs$-entry of $\mathbb{H}^{(\ell)}(\mathbb{x})]^{\sharp}$ (i.e., the square-root of $\mathbb{H}^{(\ell)}(\mathbb{x})$). Note that, given $\mathbb{H}^{(\ell-1)}(\mathbb{x})$, $\{\gamma_{j}^{(\ell-1)}(\mathbb{x})\}_{j\in[n_{\ell}]}$ are i.i.d. with law $\mathcal{N}_{d}(0,\mathbb{H}^{(\ell-1)}(\mathbb{x}))$. Note also that
@@
(A.2) $\mathbb{H}^{(\ell)}(\mathbb{x})=C_{B}1_{d}1_{d}^{\top}+C_{W}\sigma\left(\widetilde{\mathbb{N}}^{(\ell)}[\mathbb{H}^{(\ell-1)}(\mathbb{x})]^{\sharp}\right)^{\top}\mathbb{V}^{(\ell)}\sigma\left(\widetilde{\mathbb{N}}^{(\ell)}[\mathbb{H}^{(\ell-1)}(\mathbb{x})]^{\sharp}\right)$
(A.3) $\boldsymbol{\gamma}^{(\ell-1)}(\mathbb{x})=\left(\widetilde{\mathbb{N}}^{(\ell)}[\mathbb{H}^{(\ell-1)}(\mathbb{x})]^{\sharp}\right)^{\top}.$
@@
Lemma 3.
Under the foregoing assumptions and notation, it holds $\mathbb{Z}^{(L+1)}(\mathbb{x})\overset{\mathcal{L}}{=}\widetilde{\mathbb{N}}^{(L+1)}[\mathbb{H}^{(L)}(\mathbb{x})]^{\sharp}.$ Hence
(A.4) $\phi(\boldsymbol{\theta};\mathbb{Z}^{(L+1)}(\mathbb{x}))=\mathbb{E}\left[\exp\left(-\frac{1}{2}\mathrm{Tr}\big[\boldsymbol{\theta}^{\top}\mathbb{H}^{(L)}(\mathbb{x})\boldsymbol{\theta}\big]\right)\right],\quad\boldsymbol{\theta}\in\mathbb{R}^{d\times n_{L+1}}.$
Lemma 4.
Let $\{V_{n,j}\}_{n\geq 1,j\in[n]}$ be a sequence of non-negative r.v.’s such that $\sum_{j=1}^{n}V_{n,j}\overset{\mathcal{L}}{\to}{\rm ID}(a,\rho),$ as $n\to\infty$, where $a\geq 0$ is a non-negative constant and $\rho$ is a Lévy measure on $(0,\infty)$, and let $\{\xi_{j}\}_{j\geq 1}$ be a sequence of i.i.d. r.v.’s with law $\mathcal{N}_{d}(0,\mathbb{C})$, independent of $\{V_{n,j}\}_{n\geq 1,j\in[n]}$. If $\sigma$ satisfies the assumption $(i)$ of Theorem 1, then
@@
Before proving Theorem 2, we need more details about sequential limits performed in a given order. For $\ell\in\mathbb{N}^{*}$, let $f:\mathbb{N}^{\ell}\rightarrow\mathbb{R}$ be a function, and, for $S:=\{i_{1},\ldots,i_{r}\}\subseteq[\ell]$, $i_{1}<i_{2}<\ldots<i_{r}$, let $\beta:S\rightarrow[\ell]$ be an injection. By $\beta\lim f$ we mean $\lim_{n_{\beta(i_{r})}\to\infty}\lim_{n_{\beta(i_{r-1})}\to\infty}\cdots\lim_{n_{\beta(i_{1})}\to\infty}f$ and we say that the limit is sequential along $\beta$. Let $\pi$ be a permutation of $\ell$] and $j\in\{0\}\cup[\ell]$. By $(\pi\preceq j)$ we denote the injection obtained by restricting $\pi$ on $[j]$. Analogously, $(\pi\succeq j)$ denotes the injection obtained by restricting $\pi$ on $\{j,j+1,\ldots,\ell\}$. If $j=0$, we denote by ($\pi\preceq 0$) the unique injection $\emptyset\to[L]$ and we set $(\pi\preceq 0)\lim f:=f$. Clearly, if $j=\ell$, then $(\pi\preceq j)=\pi$ while $(\pi\succeq j)$ is the unique permutation of $\{\ell\}$. Moreover, if $j\not$$\ell$, then
(A.5) $\pi\lim f=(\pi\succeq j+1)\lim\left[(\pi\preceq j)\lim f\right].$
The preceding definitions can be extended in a natural way to other notions of limits. For instance, if $f$ is random, we write $f\xrightarrow{\beta\mathcal{L}}\overline{f}$ to mean that $f$ converges in law to $\overline{f}$ sequentially along $\beta$. Analogously, we write $f\xrightarrow{\beta\mathbb{P}}\overline{f}$ to mean that $f$ converges to $\overline{f}$ $\mathbb{P}$-a.s. sequentially along $\beta$. Hereon, if $\pi$ is a permutation of $\ell$], and $f\xrightarrow{\pi\mathcal{L}}\overline{f}$, then we denote by $\overline{f}_{(\pi\preceq j)}$ the weak sequential limit of $f$ along $(\pi\preceq j)$.
Proof of Theorem 2.
For ease of notation, we omit the explicit reference to the input matrix $\mathbb{x}$ in every symbol it occurs. Since vec and $\top$ are isometries of Euclidean spaces, we equivalently prove the theorem for the random matrices $\mathbb{Z}^{(L+1)}$ and $\mathbb{G}^{(L+1)}$ rather that for the r.v.’s $Z^{(L+1)}$ ad $G^{(L+1)}$. Let $\mathcal{S}_{+}^{d}$ be the space of the order $d$ symmetric and positive semi-definite real matrices endowed with the Frobenius norm, and note that, for any fixed $\boldsymbol{\theta}\in\mathbb{R}^{d\times d}$, the mapping $\mathcal{S}_{+}^{d}\ni\mathbb{A}\mapsto\exp\left(-\frac{1}{2}\mathrm{Tr}\big[\boldsymbol{\theta}^{\top}\mathbb{A}\boldsymbol{\theta}\big]\right)$ is real-valued, continuous and bounded. Hence, by (A.4) and the definition of weak convergence it holds
$\mathbb{H}^{(L)}\xrightarrow{\pi\mathcal{L}}\mathbb{K}^{(L+1)}\Longrightarrow\pi\lim\phi(\boldsymbol{\theta};\mathbb{Z}^{(L+1)})=\mathbb{E}\left[\exp\left(-\frac{1}{2}\mathrm{Tr}\big[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}\boldsymbol{\theta}\big]\right)\right]=\phi(\boldsymbol{\theta};\mathbb{G}^{(L+1)}).$
Therefore to prove the theorem it suffices to prove that $\mathbb{H}^{(L)}\xrightarrow{\pi\mathcal{L}}\mathbb{K}^{(L+1)}$, for every permutation $\pi$ of $[L]$. We proceed by induction on the number of layers $L$. The basis of the induction: $\mathbb{H}^{(1)}\xrightarrow{\pi\mathcal{L}}\mathbb{K}^{(2)}$, reduces to
(A.6) $\mathbb{H}^{(1)}{\xrightarrow[n_{1}\to\infty]{\mathcal{L}}}\mathbb{K}^{(2)}$
since, for $L=1$, the unique permutation of $[L]$ is the identity. We defer the proof of (A.6) after the proof of the inductive step. So, let $L\geq 3$ and assume
(A.7) $\mathbb{H}^{(L-1)}\xrightarrow{\tau\mathcal{L}}\mathbb{K}^{(L)}\quad\text{for every permutation $\tau$ of $[L-1]$}.$
@@
(A.8) $\pi\lim\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})=\phi(\boldsymbol{\theta};\mathbb{K}^{(L+1)})\quad\text{$\forall\boldsymbol{\theta}\in\mathbb{R}^{d\times d}$ and every permutation $\pi$ of $[L]$},$
assuming (A.6) (to be proved) and (A.7). Let $\pi$ be any permutation of $[L]$ and let $j^{*}=\pi^{-1}(L)$. Hence $(\pi\preceq j^{*}-1)=(\tau\preceq j^{*}-1)$ for some permutation $\tau$ of [L1][L-1]. Since the mapping $\mathbb{A}\to\mathbb{A}^{\sharp}$ is continuous on $\mathcal{S}_{+}^{d}$, the inductive step (A.7) yields $\mathbb{H}^{(L-1)}]^{\sharp}\xrightarrow{(\pi\preceq j^{*}-1)\mathcal{L}}\mathbb{K}^{(L)}_{(\pi\preceq j^{*}-1)}$. Since, by construction, $\mathbb{H}^{(L-1)}$ is independent of $\mathbb{V}^{(L)}$ and $\widetilde{\mathbb{N}}^{(L)}$ and $\sigma$ is continuous, by (A.2) it follows that
$\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{H}^{(L-1)}]^{\sharp}\right)^{\top}\!\mathbb{V}^{(L)}\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{H}^{(\ell-1)}]^{\sharp}\right)$
converges to
$\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}\right)^{\top}\mathbb{V}^{(L)}\sigma\left(\widetilde{\mathbb{N}}^{(L)}[\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}\right)$
in law along $(\pi\preceq j^{*}\!-\!1)$}. Note that also $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}-1)}]^{\sharp}$ is independent of $\mathbb{V}^{(L)}$ and $\widetilde{\mathbb{N}}^{(L)}$. Therefore, if $\mathbb{\Xi}:=[\xi_{1},\ldots,\xi_{n_{L}}]\in\mathbb{R}^{d\times n}$ is the random matrix defined by $\mathbb{\Xi}=\widetilde{\mathbb{N}}^{(L)}[\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}$, then $\mathbb{\Xi}$ is independent of $\mathbb{V}^{(L)}$ and, given $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]^{\sharp}$, $\xi_{1},\ldots,\xi_{n_{L}}$ are i.i.d. $d$-dimensional r.v.’s with law $\mathcal{N}_{d}(0,\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)})$. Consequently, in view of (A.1), one has
$(\pi\preceq\!j^{*}\!-\!1)\lim\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})=\mathbb{E}\left[\exp\left(\mathbb{i}\langle\boldsymbol{\theta},C_{B}\mathbb{1}_{d}\mathbb{1}_{d}^{\top}+C_{W}\sum_{j=1}^{n_{L}}V_{n_{L}j}^{(L)}\sigma(\xi_{j})\sigma(\xi_{j})^{\top}\rangle_{F}\right)\right].$
On the other hand, with $n_{\pi(j^{*})}=n_{L}$, by Lemma 4 we have
@@
where, referring to (4.3), $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}$ is defined similarly to $\mathbb{K}^{(L+1)}$ but with the $\xi_{j}$’s in place of the $\zeta_{j}^{(L)}$’s and $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}$ in place of $\mathbb{K}^{(L)}$, i.e.,
(A.9) $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}:=C_{B}1_{d}1_{d}^{\top}+C_{W}\Biggl(a^{(L)}\mathbb{E}[\sigma(\xi_{1})\sigma(\xi_{1})^{\top}\,|\,\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}]+\sum_{j=1}^{N_{L}((0,\infty))}T_{j}^{(L)}\sigma(\xi_{j})\sigma(\xi_{j})^{\top}\Biggr).$
Since $L=\pi(j^{*})$ we have proved
(A.10) $(\pi\preceq j^{*})\lim\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})=\lim_{n_{L}\to\infty}\left[(\pi\preceq\!j^{*}\!-\!1)\lim\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})\right]=\mathbb{E}\left[\exp\left(\mathbb{i}\mathrm{Tr}\left[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}\right]\right)\right].$
If $j^{*}=L$, then we are done. Indeed, since $(\pi\preceq j^{*})=\pi$ and $(\pi\preceq j^{*}\!-\!1)=\tau$ for some permutation $\tau$ of [L1][L-1], it follows that $\mathbb{K}^{(L)}_{(\pi\preceq j^{*}\!-\!1)}=\mathbb{K}^{(L)}$ by (A.7). Hence $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}=\mathbb{K}^{(L+1)}$ by (A.9), after recalling (4.3). We thus assume $j^{*}\not=L$ and prove that $\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}\xrightarrow{(\pi\succeq j^{*}+1)\mathcal{L}}\mathbb{K}^{(L+1)}$. Since $(\pi\succeq j^{*}\!+\!1)=(\tau\succeq j^{*}\!+\!1)$ for some permutation $\tau$ of $[L]$, it follows that $\mathbb{K}^{(L)}_{(\pi\preceq j^{*})}\xrightarrow{(\pi\succeq j^{*}+1)\mathcal{L}}\mathbb{K}^{(L)}$ by (A.7) . A simple computation with characteristic functions shows
(A.11) $(\xi_{1},\ldots,\xi_{n_{L}})\xrightarrow{(\pi\succeq j^{*}+1)\mathcal{L}}(\zeta_{1}^{(L)},\ldots,\zeta_{n_{L}}^{(L)}).$
@@
$\tilde{\mathbb{E}}\left[\sigma(\xi_{1})\sigma(\xi_{1})^{\top}\,|\,\mathbb{K}^{(L)}_{(\pi\preceq j^{*})}\right]\xrightarrow{(\pi\succeq j^{*}+1)\tilde{\mathbb{P}}}\tilde{\mathbb{E}}\left[\sigma(\zeta_{1}^{(L)}\sigma(\zeta_{1}^{(L)})^{\top}\,|\,\mathbb{K}^{(L)}\right].$
Using the continuity of $\sigma$, one also has
$\sum_{j=1}^{N_{L}((0,\infty))}T_{j}^{(L)}\sigma(\xi_{j})\sigma(\xi_{j})^{\top}\xrightarrow{(\pi\succeq j^{*}+1)\tilde{\mathbb{P}}}\sum_{j=1}^{N_{L}((0,\infty))}T_{j}^{(L)}\sigma(\zeta_{j}^{(L)})\sigma(\zeta_{j}^{(L)})^{\top}$
@@
$\begin{split}\pi\lim\phi(\boldsymbol{\theta};\mathbb{H}^{(L)})&=(\pi\succeq j^{*}+1)\lim\left[(\pi\preceq j^{*})\lim\phi(\boldsymbol{\theta};\mathbb{H}^{L})\right]\\ &=(\pi\succeq j^{*}+1)\mathbb{E}\left[\exp\left(\mathbb{i}\mathrm{Tr}\left[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}_{(\pi\preceq j^{*})}\right]\right)\right]=\mathbb{E}\left[\exp\left(\mathbb{i}\mathrm{Tr}\left[\boldsymbol{\theta}^{\top}\mathbb{K}^{(L+1)}\right]\right)\right]\end{split}$
and (A.8) is proved. It remains to prove (A.6). By construction $\mathbb{H}^{(0)}=\mathbb{K}^{(1)}$. Hence, with $\ell=L=1$, by (A.3), for $j\in[n_{1}]$, it holds that $\gamma_{j}^{(0)}\overset{\mathcal{L}}{=}\zeta_{j}^{(1)}$, where $\zeta_{j}^{(1)}\sim\mathcal{N}_{d}(0,\mathbb{K}^{(1)})$. Consequently, by (A.1), one has
@@
Hence, after recalling (4.3), by Lemma 4, $\mathbb{K}^{(2)}$ is the limit in law of $\mathbb{H}^{(1)}$ as $n_{1}\to\infty$. The proof is thus completed. ∎